find the domain
8(x)=5/1-3x

Answer by stanbon(75887)

find the domain
8(x)=5/1-3x
-----
Your notation is confusing.
----
But it looks like 1-3x cannot be zero
Solve: 1-3x = 0
x = 1/3
----------------
Domain: All Real numbers except x = 1/3
========================
Cheers,
Stan H.
========================